At a given point above the surface of Earth, the gravitational acceleration is equal to 7.3 m/s2. What is the altitude of this point, above the surface of Earth, in km.
Answer #1

here,

mass of earth , M = 5.98 * 10^24 kg

radius , r = 6.371 * 10^6 m

let the altitude above the earth surface be A

the gravity acceleration , a = G * m /(r + A)^2

7.3 = 6.67 * 10^-11 * 5.98 * 10^24 /(6.371 * 10^6 + A)^2

solving for A

A = 1.02 * 10^6 m = 1020 km

the altitude above the earth surface is 1020 km